# Stable Collisionless Tori Around Kerr Black Holes

### v1

@@ 5.1 Bulk Motion
It turns out that $\Omega$ is the same in both the Boyer-Lindquist coordinates and the Kerr-Schild coordinates. Figure 2 shows the fluid angular velocity $\Omega$ in the equatorial plane. It can be seen that $\Omega$ is super-Keplerian at small radii but sub-Keplerian at large radii. This is similar to the Fishbone-Moncrief torus (L. G. Fishbone & V. Moncrief, 1976). Due to the differential rotation, we expect this torus to be an excellent basis for studying the magnetorotational instability (MRI) in the accretion flow.
@@ 5.2 Effective Temperature
The GRPIC framework allows us to track individual particle trajectories and compute thermodynamic quantities from first principles. Since temperature is only well defined in the fluid rest frame (FRF), we first transform particle velocities accordingly. It is most convenient to construct the FRF tetrad from the Boyer-Lindquist coordinates, as we show in Appendix E. Therefore, to obtain particle velocities in the FRF, we first transform their Kerr-Schild 4-velocity, $v_{\mu}^{\rm(KS)}$, to Boyer-Lindquist coordinates, $v_{\mu}^{\rm(BL)}$, then apply the transformation (E4) to get the FRF orthonormal 4-velocity $v_{\hat{\mu}}$. The effective temperature is defined as
@@
Panel I of Figure 3 shows how $T_{\mathrm{eff}}$ changes over the torus, and panel II shows the probability density of $v_{\mathrm{FRF}}$ at different places on the torus. For this calculation we used a special run with $1000$ particles per cell to ensure excellent statistics, and the physical parameters are identical to the second row of Figure 1 with $r_{\rm in}=10\,r_{g}$. It can be seen that near the point $(r=r_{0},\theta=\pi/2)$ where $E_{\max}-E_{\min}$ takes the largest range and $T<E_{\max}-E_{\min}$, the distribution is more or less Maxwellian. Toward the edge of the torus where $E_{\min}$ approaches $E_{\max}$, the distribution of $v_{\rm FRF}$ is progressively truncated from the higher end, and the effective temperature goes to zero as expected.
@@ 5.3 Pressure Tensor
This expression can be evaluated analytically for the equilibrium distribution, though the resulting form is not compact. The symmetries of the problem simplify the structure of $T_{\mu\nu}$. In Boyer-Lindquist coordinates, we find:
@@
Given the large pressure anisotropy, the plasma should be unstable under the Weibel instability (see e.g. N. A. Krall & A. W. Trivelpiece, 1973). However, in our 2D simulations, the $\phi$ direction is taken to be the symmetry direction, which suppresses the unstable modes since their wave vectors are along the colder direction. At the same time, differential rotation may also introduce large-scale phase mixing which has been demonstrated to spontaneously produce seed magnetic fields (M. Zhou et al., 2022). Here we are again protected from this type of instability since the shearing flow direction is in our symmetry direction. In a full 3D simulation, we expect that Weibel modes will grow in this setup and spontaneously generate magnetic fields, which will eventually destabilize the torus.
@@ 6 Discussion
In this paper, we have constructed a fully analytic equilibrium torus solution with finite angular momentum in the collisionless regime around a Kerr black hole. We realized this torus solution in a 2D axisymmetric GRPIC code and verified that it is indeed stable over hundreds of dynamic times under zero electromagnetic field. By adding a seed magnetic field, this initial condition will become unstable to the magnetorotational instability (MRI), shedding angular momentum and start the accretion process. Therefore, this solution can serve as an ideal initial condition for studying collisionless accretion in a GRPIC framework, allowing the controlled development of MRI and angular momentum transfer, similar to the Fishbone-Moncrief torus widely used in GRMHD simulations. A host of open problems regarding this type of collisionless accretion flow around low luminosity AGN can now be addressed.
@@
The strong pressure anisotropy in the initial condition may lead to the development of the Weibel instability. 2D axisymmetric simulations seem to be protected from this instability, while it may play a role in full 3D setups. If this feature is undesirable, one should be able to finetune the profile of $L_{z}$ such that the $\phi$ temperature in the fluid rest frame matches the temperature in the other two directions. On the other hand, if sufficient seed magnetic field is put in initially, then it can also suppress the growth of the Weibel instability, allowing the system to develop MRI self-consistently.
@@ Appendix A Metric
The Kerr spacetime can be parameterized in Boyer-Lindquist coordinates
@@
where $\Delta=r^{2}-2Mr+a^{2}$, $\Sigma=r^{2}+a^{2}\cos^{2}\theta$ and $A=(r^{2}+a^{2})^{2}-a^{2}\sin^{2}\theta\Delta$. Using the transformation given by e.g.  S. S. Komissarov (2004), we can convert a vector $V^{\mu}$ from Boyer-Lindquist into Kerr-Schild $V^{\mu}$$^{\prime}$
@@
We can write the Kerr spacetime in terms of Kerr-Schild coordinates
@@
Noting that the spatial metric $\gamma_{ij}$ is used to raise and lower spatial vectors e.g. $\beta_{i}=\gamma_{ij}\beta^{j}$. We assume the metric is stationary $\partial_{t}g=0$ which implies $\alpha$, $\beta^{i}$ and $\gamma_{\mu\nu}$ are also time independent. Explicitly these the Boyer-Lindquist components are (spatial components are ordered by $r$, $\theta$, $\phi$):
@@ Appendix C Separation of Variables
Subtracting $2aEL$ from both sides
@@ Appendix E Tetrad of the fluid rest frame
To evaluate the fluid properties of the torus, we must compute quantities in the rest frame of the bulk flow, the fluid rest frame (FRF). Here we construct the transformation from Boyer-Lindquist coordinates directly into the FRF. In the FRF, the bulk four velocity of the fluid must be $u^{\hat{\mu}}=(1,0,0,0)$. Thus the transformation $e^{\mu}_{\hskip 2.45836pt{\hat{\nu}}}$ must take $u^{\hat{\mu}}\to u^{\mu}$. So
@@
This constrains the zeroth basis vector to $e_{\hat{0}}=u^{0}\partial_{t}+u^{\phi}\partial_{\phi}$. Since Boyer-Lindquist coordinates are diagonal for $r,\theta$, we choose $e_{\hat{r}}=g_{rr}^{-1/2}\partial_{r}$ and $e_{\hat{\theta}}=g_{\theta\theta}^{-1/2}\partial_{\theta}$. This structure maintains the direction of the $r$ and $\theta$ components. We choose the final basis vector to be orthogonal to the rest. This has the form $e_{\hat{\phi}}=c_{1}\partial_{t}+c_{2}\partial_{\phi}$. The orthogonality ensures that
@@ Appendix F Stress Tensor
In Section 5.3, we described that the components of the stress energy tensor can be integrated analytically in Boyer-Lindquist coordinates. To evaluate these integrals, we expand (29) using the coordinate transform (14) as

### v2

@@ 5.1 Bulk Motion
$\Omega$ is the same in both the Boyer–Lindquist coordinates and the Kerr–Schild coordinates. Figure 2 shows the fluid angular velocity $\Omega$ in the equatorial plane. It can be seen that $\Omega$ is super-Keplerian at small radii but sub-Keplerian at large radii. This is similar to the Fishbone-Moncrief torus (L. G. Fishbone & V. Moncrief, 1976). Due to the differential rotation, we expect this torus to be an excellent basis for studying the magnetorotational instability (MRI) in the accretion flow.
@@ 5.2 Effective Temperature
The GRPIC framework allows us to track individual particle trajectories and compute thermodynamic quantities from first principles. Since temperature is only well defined in the fluid rest frame (FRF), we first transform particle velocities accordingly. It is most convenient to construct the FRF tetrad from the Boyer–Lindquist coordinates, as we show in Appendix E. Therefore, to obtain particle velocities in the FRF, we first transform their Kerr–Schild 4-velocity, $v_{\mu}^{\rm(KS)}$, to Boyer–Lindquist coordinates, $v_{\mu}^{\rm(BL)}$, then apply the transformation (E4) to get the FRF orthonormal 4-velocity $v_{\hat{\mu}}$. The effective temperature is defined as
@@
Panel I of Figure 3 shows how $T_{\mathrm{eff}}$ changes over the torus, and panel II shows the probability density of $v_{\mathrm{FRF}}$ at different places on the torus. For this calculation we used a special run with $1000$ particles per cell to ensure excellent statistics, and the physical parameters are identical to the second row of Figure 1 with $r_{\rm in}=10\,r_{g}$. It can be seen that near the point $(r=r_{0},\theta=\pi/2)$ where $E_{\max}-E_{\min}$ is largest and $T<E_{\max}-E_{\min}$, the distribution is more or less Maxwellian. Toward the edge of the torus where $E_{\min}$ approaches $E_{\max}$, the distribution of $v_{\rm FRF}$ is progressively truncated from the higher end, and the effective temperature goes to zero as expected.
@@ 5.3 Pressure Tensor
This expression can be evaluated analytically for the equilibrium distribution, though the resulting form is not compact. The symmetries of the problem simplify the structure of $T_{\mu\nu}$. In Boyer–Lindquist coordinates, we find:
@@
Given the large pressure anisotropy, the plasma is prone to the Weibel instability (see e.g. N. A. Krall & A. W. Trivelpiece, 1973). However, in our 2D simulations, axisymmetry suppresses the unstable modes since their wave vectors are along the colder direction. At the same time, differential rotation may also introduce large-scale phase mixing which has been demonstrated to spontaneously produce seed magnetic fields (M. Zhou et al., 2022). Here we are again protected from this type of instability since the shearing flow direction is in our symmetry direction. In a full 3D simulation, we expect that Weibel modes will grow in this setup and spontaneously generate magnetic fields, which will eventually destabilize the torus.
@@ 6 Discussion
In this paper, we have constructed a fully analytic equilibrium torus solution with finite angular momentum in the collisionless regime around a Kerr black hole. We realized this torus solution in a 2D axisymmetric GRPIC code and verified that it is indeed stable over hundreds of dynamical times under zero electromagnetic field. By adding a seed magnetic field, this initial condition will become unstable to the magnetorotational instability (MRI), shedding angular momentum and start the accretion process. Therefore, this solution can serve as an ideal initial condition for studying collisionless accretion in a GRPIC framework, allowing the controlled development of MRI and angular momentum transfer, similar to the Fishbone-Moncrief torus widely used in GRMHD simulations. A host of open problems regarding this type of collisionless accretion flow around low-luminosity AGN can now be addressed.
@@
The strong pressure anisotropy in the initial condition leads to the development of the Weibel instability. 2D axisymmetric simulations suppress this instability, but the instability develops in full 3D setups. If this feature is undesirable, one should be able to finetune the profile of $L_{z}$ such that the $\phi$ temperature in the fluid rest frame matches the temperature in the other two directions. On the other hand, if sufficient seed magnetic field is put in initially, then it can also suppress the growth of the Weibel instability, allowing the system to develop MRI self-consistently.
We have briefly experimented with adding a weak magnetic field to the initial condition in a full PIC simulation. When the magnetic field is sufficiently weak (e.g. gyroradius much larger than the size of the torus), the torus remains stable for a long time, regardless of the field configuration. With a larger magnetic field such that the plasma is magnetized, gyration tends to equilibrate the pressure along the two directions perpendicular to the magnetic field. As a result, the strong initial pressure anisotropy in our equilibrium torus quickly develops into an anisotropy with respect to the magnetic field direction, $p_{\perp}<p_{\parallel}$. In the center of the torus where $\beta\gg 1$, this leads to the development of the firehose instability (see e.g. M. W. Kunz et al., 2014) that brings the plasma closer to isotropy. Long term evolution of this scenario invariably leads to the development of MRI which initiates the accretion of the torus onto the black hole. We will explore the physics of collisionless accretion in a future work.
@@ Appendix A Metric
The Kerr spacetime can be parameterized in Boyer–Lindquist coordinates
@@
where $\Delta=r^{2}-2Mr+a^{2}$, $\Sigma=r^{2}+a^{2}\cos^{2}\theta$ and $A=(r^{2}+a^{2})^{2}-a^{2}\sin^{2}\theta\Delta$. Using the transformation given by e.g.,  S. S. Komissarov (2004), we can convert a vector $V^{\mu}$ from Boyer–Lindquist into Kerr–Schild $V^{\mu}$$^{\prime}$
@@
We can write the Kerr spacetime in terms of Kerr–Schild coordinates
@@
Noting that the spatial metric $\gamma_{ij}$ is used to raise and lower spatial vectors e.g. $\beta_{i}=\gamma_{ij}\beta^{j}$. We assume the metric is stationary $\partial_{t}g=0$ which implies $\alpha$, $\beta^{i}$ and $\gamma_{\mu\nu}$ are also time independent. Explicitly these the Boyer–Lindquist components are (spatial components are ordered by $r$, $\theta$, $\phi$):
@@ Appendix C Separation of Variables
Subtracting $2aEL_{z}$ from both sides
@@ Appendix E Tetrad of the fluid rest frame
To evaluate the fluid properties of the torus, we must compute quantities in the rest frame of the bulk flow, the fluid rest frame (FRF). Here we construct the transformation from Boyer–Lindquist coordinates directly into the FRF. In the FRF, the bulk four velocity of the fluid must be $u^{\hat{\mu}}=(1,0,0,0)$. Thus the transformation $e^{\mu}_{\hskip 2.45836pt{\hat{\nu}}}$ must take $u^{\hat{\mu}}\to u^{\mu}$. So
@@
This constrains the zeroth basis vector to $e_{\hat{0}}=u^{0}\partial_{t}+u^{\phi}\partial_{\phi}$. Since Boyer–Lindquist coordinates are diagonal for $r,\theta$, we choose $e_{\hat{r}}=g_{rr}^{-1/2}\partial_{r}$ and $e_{\hat{\theta}}=g_{\theta\theta}^{-1/2}\partial_{\theta}$. This structure maintains the direction of the $r$ and $\theta$ components. We choose the final basis vector to be orthogonal to the rest. This has the form $e_{\hat{\phi}}=c_{1}\partial_{t}+c_{2}\partial_{\phi}$. The orthogonality ensures that
@@ Appendix F Stress Tensor
In Section 5.3, we described that the components of the stress energy tensor can be integrated analytically in Boyer–Lindquist coordinates. To evaluate these integrals, we expand (29) using the coordinate transform (14) as